# DiSCo: Device-Server Collaborative LLM-Based Text Streaming Services

### v1

@@ 1 Introduction
Large language models (LLMs) have revolutionized various applications, with over 60% focusing on conversational interactions such as chatbots (Grand View Research, 2023). Meeting high serving demands requires scaling deployments across on-premise servers in the cloud and on-device inference, as seen in Apple Intelligence (Gunter et al., 2024) and Google’s Gemini Nano (Google, 2024). The Quality of Experience (QoE) for interactive applications is primarily evaluated by two critical metrics: Time-To-First-Token (TTFT) in the prefill stage, which quantifies the initial response latency, and Time-Between-Token (TBT) during the decode stage, which measures the consistency of token delivery speed (Databricks, 2023; Liu et al., 2024a, b).
On-server deployments lower serving costs by sharing infrastructure among many requests but often introduce unpredictable high latency due to request queueing delays (Agrawal et al., 2024), and the internet speed to end users fluctuates. While on-device deployment enables increasingly capable LLMs with sufficient accuracy, it suffers from slow processing speeds for long prompts and high energy consumption. For example, an iPhone running a 7B parameter LLM can operate for less than two hours on a full charge (Liu et al., 2024c).
This paper introduces a novel paradigm for cost-constrained device-server cooperative inference. We incorporate both server usage (e.g., monetary costs) and device energy costs via a dynamic exchange rate, which end users can adjust to balance response generation between the cloud and devices. As such, we can strategically distribute inference requests between endpoints and dynamically migrate ongoing token generation to maximize QoE. However, realizing this vision presents several fundamental challenges:
  • Unified Cost Management: The total serving cost combines heterogeneous resource expenditures from both endpoints—monetary costs from server API usage and energy costs from device computation. The relative value of energy costs varies dynamically based on device context (e.g., battery level, charging status) and user preferences for server spending, making it challenging to establish a unified optimization.
@@
  • Migration Impact on Token Delivery: While dynamic migration between endpoints can reduce overall running costs, it risks disrupting TBT. The challenge lies in determining when and how to perform migration while minimizing the degradation of user experience and the increase in costs.
@@
  • Cost-Aware Dispatching Policies: We introduce two dispatching mechanisms targeting different cost constraints. For server cost constraints, we employ a length-threshold based dispatching that routes requests shorter than a dynamically computed threshold to devices. For device energy constraints, we implement a delay-based dispatching mechanism where devices wait for a computed interval before starting local inference. Both mechanisms adapt their thresholds based on unified cost measures that combine server monetary costs and device energy consumption.
  • Token-Level Migration Framework: We enable seamless generation handoff between endpoints through a novel migration protocol that preserves token delivery consistency. Our framework employs delayed migration timing to minimize interruption, while a token buffer ensures smooth delivery during transitions. This design maintains user experience while saving resource costs across endpoints.
Through extensive evaluation using real-world traces from commercial LLM streaming API services (including GPT and DeepSeek) and on-device deployments, we demonstrate that DiSCo improves mean and tail TTFT by up to 50% without violation of TBT, significantly reducing the costs.
@@ 2.1 LLM Token Mixture and Routing
Device-server collaborative approaches have evolved along two directions. First, systems like EdgeShard Zhang et al. (2024) and WDMoE (Xue et al., 2024a) partition LLMs across multiple endpoints when a single device cannot host the entire model. LLMCad Xu et al. (2023) uses on-device models to reduce server costs, while PerLLM (Yang et al., 2024) optimizes energy consumption across devices and servers under constraints. Second, routing-based approaches (Ong et al., 2024; Ding et al., 2024) balance cost and accuracy by directing simple requests to small models and complex queries to advanced ones. However, these approaches do not optimize token delivery metrics (TTFT and TBT) under cost constraints.
@@ 2.2 LLM-Based Text Streaming Applications
Current LLM systems struggle to meet user expectations for these metrics, with TTFTs ranging from hundreds of milliseconds to over ten seconds—far exceeding the ideal latencies of tens of milliseconds for interactive applications Mäki-Patola and Hämäläinen (2004); Žádník et al. (2022). Token consumption patterns vary by output modality: in visual text scenarios, reading speeds differ across demographic groups, with the majority (52%) aged 25-44 reading at 4-5 tokens per second, while older groups generally read more slowly Liu et al. (2024a); Brysbaert (2019); Petrov et al. (2024). Audio output consumption shows more consistency, averaging 3-4 tokens per second across languages Liu et al. (2024a); Parachuk (2022); Barnard, Dom (2022). Notably, conventional evaluation metrics like token generation throughput or average time-per-output-token provide incomplete insights, as they fail to capture the crucial relationship between token delivery timing and actual user consumption patterns.
@@ 2.3 Limitations of Existing Text Streaming Applications
Existing LLM serving relies on two deployment paradigms: on-server and on-device inference. With rapid hardware and software advancements, on-device LLMs have achieved accuracy levels sufficient for many applications, as evidenced by the integration of Apple Intelligence (Gunter et al., 2024) and Google’s Gemini Nano Google (2024) into iOS and Android platforms, where they effectively handle text completion and message composition tasks. While on-device LLMs may still be inadequate for complex tasks (e.g., advanced mathematical reasoning), we focus on the growing category of applications where current on-device models already achieve satisfactory accuracy. For these applications, the challenge is not model capability, but rather the substantial monetary or energy cost demands of LLM inference.
Unfortunately, both serving paradigms face challenges. On-device inference, despite enabling faster generation owing to its dedicated resources (Song et al., 2023; Xue et al., 2024b), suffers from extended TTFT for long prompts due to limited processing speeds, and substantial energy consumption that scales linearly with response lengths (Li et al., 2024b). For instance, a fully charged iPhone running a 7B parameter LLM can operate for less than two hours (Liu et al., 2024c)—insufficient for day-long mobile use.
@@ 3 Characterizing LLM Inference
This section characterizes LLM inference performance in on-server and on-device paradigms, which informs our design.
@@
Our measurements reveal contrasting TTFT patterns between on-device and on-server inference. As shown in Figure 2, on-device inference exhibits stable TTFTs when processing identical prompts at 60-second intervals, primarily reflecting the prefill duration due to dedicated local hardware resources. In contrast, on-server inference experiences high variations and significant tail latency, attributed to network delays, request queuing, and resource contention.
We summarize the TTFT performance of 1,000 requests in Table 1. We observe that on-device TTFT scales linearly with prompt length due to hardware constraints (Li et al., 2024b), while on-server TTFT shows minimal prompt-length sensitivity through advanced resource scaling (Zhong et al., 2024; Patel et al., 2024; Hu et al., 2024).
@@
TBT characterizes the I/O-bound decode stage latency. Analysis of temporal samples and distributions across six setups (Figure 3) reveals higher TBT variability in on-server inference compared to on-device execution. More importantly, both deployment approaches achieve generation speeds exceeding user consumption rates (§2.2), making cooperative serving practical.
@@
Our studies further reveal that as on-device models continue to improve—often fine-tuned for specific tasks (Gunter et al., 2024; Liu et al., 2024c)—their performance increasingly matches that of on-server models in popular applications like instruction-following and translation (detailed in §5 and Appendix D). However, deploying these models on-device introduces challenges such as long prefilling latency and startup overhead.
@@
Taking these findings together—particularly the predictable performance of on-device inference and the elastic scaling capabilities of server-based inference—we observe opportunities for optimization in cost-constrained device-server cooperative serving. Dynamic request migration between server and device endpoints during response generation can yield significant cost savings.
@@ 4.1 Problem Formulation
We propose a unified cost model combining both monetary bills from on-server inference and energy bills from on-device inference. Let $c^{p}_{s}$ and $c^{d}_{s}$ denote the per-token monetary costs for server prefill and decode phases respectively, while $c^{p}_{d}$ and $c^{d}_{d}$ represent the per-token energy costs for device prefill and decode phases. Converting between energy and monetary costs is done by a dynamic exchange rate $\lambda$, adjusted by users to reflect their preferences. We offer a user-friendly tunable budget ratio $b\in[0,1]$, representing the additional cost allowance beyond baseline costs. Our optimization objectives focus on: (1) minimizing both mean and tail TTFT, and (2) maintaining consistent token delivery at a specified pace (i.e., stable TBT).
@@ 4.2 Dispatch Controller: Cost-Aware Request Routing
Based on our analysis in §3, server-side TTFT shows weak correlation with prompt length due to various factors (network delay, request queueing, etc.). We model server TTFTs as a known distribution, obtained either from server-provided information or device-side profiling. In contrast, device-side TTFT exhibits a linear relationship with prompt length, with the coefficient determined through offline profiling.
Our key insight is that the optimization problem naturally decomposes into two scenarios based on dominant cost factors: device-constrained scenarios where energy consumption is the primary bottleneck, and server-constrained scenarios where API monetary costs dominate. This decomposition enables efficient solutions. Pseudocode for the dispatch controller is in Appendix F.
@@
When device costs dominate ($\min(c^{p}_{d},c^{d}_{d})>\max(c^{p}_{s},c^{d}_{s})$), we need to carefully manage device resource usage under a budget constraint $\mathbb{E}[I_{d}(l)l]\leq b\cdot\mathbb{E}[l]$, where $l$ is the prompt length and $I_{d}(l)$ indicates device execution. The key challenge is balancing between two goals: leveraging device execution to bound worst-case latency while conserving energy on shorter prompts where possible.
Our solution uses a wait-time strategy: for each prompt of length $l$, we first try server execution and wait for time $w(l)$ before potentially starting device execution. This conserves device energy when the server responds quickly. We determine the optimal wait time through a two-phase approach:
  • Phase 1 (Tail Protection): We reserve budget portion $\alpha$ for worst-case scenarios by setting a maximum wait time $w_{tail}=F^{-1}(1-\min(\alpha,b))$, where $F(\cdot)$ is the server TTFT distribution. This ensures we have device resources ready when server latency exceeds its $(1-\min(\alpha,b))$-th percentile.
@@
When server costs dominate ($\max(c^{p}_{s},c^{d}_{s})>\min(c^{p}_{d},c^{d}_{d})$), we need to carefully manage server resource usage under a budget constraint $\mathbb{E}[I_{s}(l)l]\leq b\cdot\mathbb{E}[l]$, where $I_{s}(l)$ indicates server execution. Our analysis in §3 shows that device TTFT scales linearly with prompt length as $T_{d}(l)=kl+c$, while server TTFT has minimal length correlation. This suggests a length-based routing strategy: short prompts run on device to conserve server budget, while long prompts use both endpoints to minimize TTFT.
@@ 4.3 Migration Controller: Cost-Efficient Token Delivery
When endpoints share the same vocabulary, we transmit token IDs rather than complete token representations. Additionally, we avoid transferring intermediate states (e.g., attention key-value cache) for two practical reasons: (1) endpoints often employ different model architectures optimized for their respective hardware, making state transfer incompatible, and (2) intermediate state transfer would incur significant network overhead. Migration triggers when projected cost savings exceed overhead:
@@
where $t_{m}$ is the estimated migration overhead time. Migration begins only when the buffer contains enough tokens ($B$) to mask the migration latency.
As shown in Figure 4, this design enables seamless handoff: the source endpoint (Row A) continues generation until the target endpoint (Row B) is ready, ensuring uninterrupted token delivery to users despite the underlying endpoint transition.

### v2

@@ 1 Introduction
Large language models (LLMs) have revolutionized various applications, with over 60% focusing on conversational interactions such as chatbots (Grand View Research, 2023). Meeting high serving demands requires scaling deployments across on-premise servers in the cloud and on-device inference, as seen in Apple Intelligence (Gunter et al., 2024) and Google’s Gemini Nano (Google, 2024). The Quality of Experience (QoE) for interactive applications is primarily evaluated by two critical metrics: Time-To-First-Token (TTFT) in the prefill stage, which quantifies the initial response latency, and Time-Between-Token (TBT) during the decode stage, which measures the consistency of token delivery speed (Databricks, 2023; Liu et al., 2024a, c).
On-server deployments lower serving costs by sharing infrastructure among many requests but often introduce unpredictable high latency due to request queuing delays (Agrawal et al., 2024) and the internet speed fluctuations. While on-device deployment is able to serve increasingly capable LLMs with sufficient accuracy, it suffers from slow processing speeds for long prompts and high energy consumption. For example, a fully-charged iPhone can only operate for less than two hours running an LLM with 7B parameters (Liu et al., 2024d).
This paper introduces a novel paradigm for cost-constrained device-server cooperative inference. We incorporate both server usage (e.g., monetary costs) and device energy costs via a dynamic exchange rate, which can be adjusted by endpoint users to balance response generation between the cloud and devices. As such, we can strategically distribute inference requests between endpoints and dynamically migrate ongoing token generation to maximize QoE. However, realizing this vision presents several fundamental challenges:
  • Unified Cost Management: Total serving costs include resource expenditures from both endpoints—monetary costs from server API usage and energy costs from device computation. The relative value of energy costs varies dynamically based on device context (e.g., battery level, charging status) and user preferences for server spending, making it challenging to establish a unified optimization strategy.
@@
  • Migration Impact on Token Delivery: While dynamic migration between endpoints can reduce overall operating costs, it risks disrupting TBT. The challenge lies in determining when and how to migrate while minimizing user experience degradation and cost increase.
@@
  • Cost-Aware Dispatching Policies: We introduce two dispatching mechanisms targeting different cost constraints. For server cost constraints, we employ a length-threshold based dispatching mechanism that routes requests shorter than a dynamically computed threshold to devices. For device energy constraints, we implement a delay-based dispatching mechanism where devices wait for a computed interval before starting local inference. Both mechanisms adapt their thresholds based on unified cost measures that combine server monetary costs and device energy consumption.
  • Token-Level Migration Framework: We enable seamless generation handoff between endpoints through a novel migration protocol that preserves the consistency of token delivery. Our framework employs delayed migration timing to minimize interruption, while a token buffer ensures smooth delivery during transitions. This design maintains user experience while saving resource costs across endpoints.
Through extensive evaluation using real-world traces from commercial LLM streaming API services, including GPT and DeepSeek, and on-device deployments, we demonstrate that DiSCo improves mean and tail TTFT by up to 50% without TBT violation, significantly reducing costs.
@@ 2.1 LLM Token Mixture and Routing
Device-server collaborative approaches have evolved in two directions. First, systems like EdgeShard Zhang et al. (2024) and WDMoE (Xue et al., 2024a) partition LLMs across multiple endpoints when a single device cannot host the entire model. LLMCad Xu et al. (2023) uses on-device models to reduce server costs, while PerLLM (Yang et al., 2024) optimizes energy consumption across devices and servers under constraints. Second, routing-based approaches (Ong et al., 2024; Ding et al., 2024) balance cost and accuracy by directing simple requests to small models and complex queries to advanced ones. However, these approaches do not optimize token delivery metrics (TTFT and TBT) under cost constraints.
@@ 2.2 LLM-Based Text Streaming Applications
Current LLM systems struggle to meet user expectations for these metrics, with TTFTs ranging from hundreds of milliseconds to over ten seconds—far exceeding the ideal latencies of tens of milliseconds for interactive applications Mäki-Patola and Hämäläinen (2004); Žádník et al. (2022). Token consumption patterns vary by output modality: In visual text scenarios, reading speeds differ across demographic groups, with the majority (52%) aged 25-44 reading 4-5 tokens per second, while older groups generally read more slowly Liu et al. (2024a); Brysbaert (2019); Petrov et al. (2024). Audio output consumption shows more consistency, averaging 3-4 tokens per second across languages Liu et al. (2024a); Parachuk (2022); Barnard, Dom (2022). Notably, conventional evaluation metrics like token generation throughput or average time-per-output-token provide incomplete insights, as they fail to capture the crucial relationship between token delivery timing and actual user consumption patterns.
@@ 2.3 Limitations of Existing Text Streaming Applications
Existing LLM serving primarily relies on two deployment paradigms: on-device and on-server inference. With rapid hardware and software advancements, on-device LLMs have achieved sufficient accuracy levels for many applications (Phan et al., 2025), as evidenced by the integration of Apple Intelligence (Gunter et al., 2024) and Google’s Gemini Nano Google (2024) into iOS and Android platforms, where they effectively handle text completion and message composition tasks. While on-device LLMs may still be inadequate for complex tasks (e.g., advanced mathematical reasoning), we focus on the growing category of applications where current on-device models already achieve satisfactory accuracy. For these applications, the challenge is not model capability, but rather the substantial monetary or energy cost demands of LLM inference.
Unfortunately, both serving paradigms face challenges. On-device inference, though enabling faster generation powered by its dedicated resources (Song et al., 2023; Xue et al., 2024b), suffers from extended TTFT for long prompts due to limited processing speeds and substantial energy consumption that scales linearly with response lengths (Li et al., 2024b). For instance, a fully-charged iPhone running a 7B parameter LLM can only operate for less than two hours (Liu et al., 2024d)—insufficient for day-long mobile use.
@@ 3 Characterizing LLM Inference
This section characterizes the LLM inference performance of on-server and on-device paradigms, which informs our design.
@@
Our measurements reveal the contrasting TTFT patterns between on-device and on-server inference. As shown in Figure 2, on-device inference exhibits stable TTFTs when processing identical prompts at 60-second intervals, primarily reflecting the prefill duration due to dedicated local hardware resources. In contrast, on-server inference experiences high variations and significant tail latency, attributed to network delays, request queuing, and resource contention.
Unlike previous works that focus solely on prefill latency (e.g., Gim et al. (2024); Kamahori et al. (2024)) or the sum of queuing and prefill (e.g., Agrawal et al. (2024); Qin et al. (2024)), we measure user-perceived TTFT as a comprehensive sum of network, queuing, and prefill latencies. Our extensive evaluations across four major LLM providers consistently demonstrate this improvement. This methodology aligns with industry benchmarks, such as Artificial Analysis Artificial Analysis (2025)’s continuous monitoring of diverse LLM services. Notably, since most on-device LLMs struggle with prefilling or understanding long prompts, we focus on short prompts, where request queuing delay and batch completion waiting time during generation dominate the overall TTFT.
We summarize the TTFT performance of 1,000 requests in Table 1. We observe that on-device TTFT scales linearly with prompt length due to hardware constraints (Li et al., 2024b), while on-server TTFT shows minimal prompt-length sensitivity through advanced resource scaling (Zhong et al., 2024; Patel et al., 2024; Hu et al., 2024). Note that LLM providers (e.g., Microsoft Azure Microsoft Azure (2025), DeepSeek DeepSeek (2025), Together.ai Together AI (2025), Hyperbolic Labs Hyperbolic (2025)) typically do not offer explicit TTFT SLOs, likely due to the high cost and complexity of maintaining such guarantees across diverse models and prompts.
@@
TBT characterizes the I/O-bound decode stage latency. Analysis of temporal samples and distributions across six setups (Figure 3) reveals higher TBT variability in on-server inference compared to on-device execution. Moreover, both deployments achieve generation speeds exceeding user consumption rates (§ 2.2), making cooperative serving practical.
@@
Our studies further reveal that as on-device models continue to improve—often fine-tuned for specific tasks (Gunter et al., 2024; Liu et al., 2024d)—they are progressively achieving performance parity with on-server models in popular applications like instruction-following and translation (detailed in §5 and Appendix D). However, deploying these models on-device introduces challenges such as long prefill latency and startup overhead.
@@
By integrating these findings—especially the predictable performance of on-device inference and the elastic scaling capabilities of server-based inference—we observe opportunities for optimization in cost-constrained device-server cooperative serving. Dynamic request migration between server and device endpoints during response generation can yield significant cost savings.
@@ 4.1 Problem Formulation
We propose a unified cost model combining both monetary bills from on-server inference and energy bills from on-device inference. Let $c^{p}_{s}$ and $c^{d}_{s}$ denote the per-token monetary costs for server prefill and decode phases, respectively, while $c^{p}_{d}$ and $c^{d}_{d}$ represent the per-token energy costs for device prefill and decode phases. Integration of energy and monetary costs is done by a dynamic exchange rate $\lambda$, adjusted by users to reflect their preferences. We offer a user-friendly tunable budget ratio $b\in[0,1]$, representing the additional cost allowance beyond baseline costs. Our optimization objectives focus on: (1) minimizing both mean and tail TTFT, and (2) maintaining consistent token delivery at a specified pace (i.e., stable TBT).
@@ 4.2 Dispatch Controller: Cost-Aware Request Routing
Based on our analysis in §3, server-side TTFT shows weak correlation with prompt length due to various factors (network delay, request queuing, etc.). We model server TTFT as a known distribution, obtained either from server-provided information or device-side profiling. In contrast, device-side TTFT exhibits a linear relationship with prompt length, with the coefficient determined through offline profiling.
Our key insight is that the optimization problem naturally decomposes into two scenarios based on dominant cost factors: device-constrained scenarios where energy consumption is the primary bottleneck, and server-constrained scenarios where API monetary costs dominate. This decomposition enables efficient solutions. The pseudocode for the dispatch controller is attached in Appendix F.
@@
When device costs dominate ($\min(c^{p}_{d},c^{d}_{d})>\max(c^{p}_{s},c^{d}_{s})$), we need to carefully manage device resource usage under a budget constraint $\mathbb{E}[I_{d}(l)l]\leq b\cdot\mathbb{E}[l]$, where $l$ is the prompt length and $I_{d}(l)$ indicates device execution. The key challenge is balancing between two goals: leveraging device execution to bound worst-case latency while conserving energy on shorter prompts when possible.
Our solution uses a waiting-time strategy: for each prompt of length $l$, we first try server execution and wait for time $w(l)$ before potentially starting device execution. This conserves device energy when the server responds quickly. We determine the optimal wait time through a two-phase approach:
  • Phase 1 (Tail Protection): We reserve a budget portion $\alpha$ for worst-case scenarios by setting a maximum wait time $w_{tail}=F^{-1}(1-\min(\alpha,b))$, where $F(\cdot)$ is the server TTFT distribution. This ensures we have device resources ready when server latency exceeds its $(1-\min(\alpha,b))$-th percentile.
@@
When server costs dominate ($\max(c^{p}_{s},c^{d}_{s})>\min(c^{p}_{d},c^{d}_{d})$), we need to carefully manage server resource usage under a budget constraint $\mathbb{E}[I_{s}(l)l]\leq b\cdot\mathbb{E}[l]$, where $I_{s}(l)$ indicates server execution. Our analysis in §3 shows that device TTFT scales linearly with prompt length as $T_{d}(l)=kl+c$, while server TTFT has minimal length correlation. This suggests a length-based routing strategy: short prompts run on the device to conserve server budget, while long prompts use both endpoints to minimize TTFT.
@@ 4.3 Migration Controller: Cost-Efficient Token Delivery
Theoretical Migration Framework.
The token-level migration protocol ensures seamless handoffs by leveraging the gap between token generation ($r_{g}$) and consumption ($r_{c}$) rates. The migration trigger is determined by comparing the cost savings against migration overhead:
$C_{\text{migration}}=\Delta c_{\text{decode}}\cdot l_{\text{remaining}}>\text{Overhead}_{\text{migration}}$ (4)
where $\Delta c_{\text{decode}}$ is the per-token cost difference between endpoints and $l_{\text{remaining}}$ represents the expected remaining sequence length. This formulation ensures that migration only occurs when the projected cost savings exceed the overhead of transferring control between endpoints.
@@
When endpoints share the same vocabulary, we transmit token IDs rather than complete token representations. Empirical analysis using the Alpaca dataset Taori et al. (2023) with the cl100k_base tokenizer OpenAI (2022) (used by GPT-3.5/4) demonstrates that token ID transmission versus UTF-8 encoded text yields 35.62% reduction in data volume when using minimum byte encoding (3 bytes per token), and 54.40% reduction in data volume when using minimum bit encoding (17 bits per token). These efficiency gains are particularly valuable in bandwidth-constrained environments, helping to minimize latency during migration.
Additionally, we avoid transferring intermediate states (e.g., attention key-value cache) for two practical reasons: (1) endpoints often employ different model architectures optimized for their respective hardware, making state transfer incompatible, and (2) intermediate state transfer would incur significant network overhead. For models with different vocabularies, we first convert tokens to text before re-tokenizing on the target model to ensure semantic consistency.
Migration is triggered when the projected cost savings exceed overhead:
@@
where $t_{m}$ is the estimated migration overhead time. Migration begins only when the buffer contains enough tokens ($B$) to overshadow the migration latency. Importantly, the source endpoint stops generating new tokens once the buffer is filled, preventing any potential conflicts or branching during the transition. This ensures deterministic token delivery as the target endpoint takes over generation.
As shown in Figure 4, this design enables seamless handoff: the source endpoint (Row A) continues generation until the buffer is filled, then stops to allow the target endpoint (Row B) to take over, ensuring uninterrupted token delivery to users despite the underlying endpoint transition.